OpenStudy (karlsbahs):

A car accelerates at 2 meters per second per second. Assuming the car starts from rest, how much time does it need to accelerate to a speed of 30 m/s?

OpenStudy (cwrw238):

use the constant acceleration equation v = u + at where u = initial velocity, a = acceleration , v = velocity at time t and t = time

OpenStudy (cwrw238):

so replace v with, 30 , u = 0 and a = 2 and find t

OpenStudy (cwrw238):

30 = 0 + 2t find t

OpenStudy (karlsbahs):

15 = t, Thank you very much.